# Entropic Trade-off Relations in Stochastic Thermodynamics via Replica Markov Processes

### v2

@@ II.9 Entropic bound on quantum trajectory observable
We aim to quantify the uncertainty associated with quantum trajectory observables. To achieve this, we analyze the probability distribution of $N_{Q\bullet}(\Gamma_{Q})$. We derive an upper bound for the Tsallis entropy of this distribution, denoted as $P(N_{Q\bullet})$. Following the classical derivation detailed in Appendix A, we obtain
$\left|\frac{\partial}{\partial\tau}H_{q}^{\mathrm{T}}\left[P\left(N_{Q\bullet}\right)\right]\right|\leq\frac{\sqrt{q\mathcal{B}(\tau)}}{2\tau(q-1)},$ (54)
where $q=2,3,\cdots$. Equation (54) holds for arbitrary observable $N_{Q\bullet}(\Gamma_{Q})$ of the quantum trajectory $\Gamma_{Q}$. Again, for $N_{Q\circ}$, which additionally satisfies the condition of Eq. (47), the following relation holds:
@@
where $q=2,3,\cdots$ and the following condition should be satisfied:
@@
Equations (54) and (55) are quantum generalizations of Eqs. (24) and (25), respectively, and thus their derivations are the same as those of Eqs. (24) and (25). From Eq. (52), we obtain the following relation (see Appendix E for details):
@@
where $q=2,3,\cdots$. Equation (57) is a quantum generalization of Eq. (27). Equations (55) and (57) are similar in that the left-hand side is identical and their right-hand sides comprise the quantum dynamical activity. It can be shown that Eq. (57) provides a tighter bound. In the classical case, the left-hand side of Eq. (28) is the Tsallis entropy of the observable $N_{\circ}$ and the right-hand side depends on the dynamical activity $\mathfrak{a}(t=0)$ at time $t=0$. On the other hand, in the quantum case, the right-hand side of Eq. (57) depends on the time integration of the quantum dynamical activity $\mathcal{B}(t)$.
Figure 3: Illustration of the dependence of the Tsallis entropic bound for state distribution [Eq. (35)]. The Tsallis entropic bound shown in Eq. (35) depends on $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$, where $B_{\mu_{0}}$ is the initial state, the local escape rate from $B_{\mu_{0}}$. The local escape rate from $B_{\mu_{0}}$ is the sum of transition rates to which the nodes reachable from the initial state with one jump. Suppose the initial state is $B_{1}$, which is colored with dark red. Then, the reachable states are represented by purple nodes. Therefore, the local escape rate becomes $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}=W_{21}+W_{31}+W_{41}$.
@@ III Numerical simulation
We perform simulations to numerically validate the bounds for the entropy.
First, we perform numerical calculations for the bound for the trajectory observable [Eq. (28)]. We generate a continuous time Markov process, where the transition rate $W_{\mu^{\prime}\mu}$ and the initial probability distribution $p_{\mu}(0)$ are determined randomly. We consider the counting observable $N(\Gamma)$ [Eq. (5)], which simply counts the number of total jumps within the interval $0,\tau]$, where $\tau$ is determined randomly. The range for these random parameters is shown in the caption of Fig. 4. We repeat this procedure to obtain random realizations and calculate the Tsallis entropy and the dynamical activity. For each random realization, we plot $H_{q}^{\mathrm{T}}[P(N_{\circ})]$ versus $\left(1-e^{-q\tau\mathfrak{a}(t=0)}\right)/(q-1)$, which are the left and right-hand sides of Eq. (28), respectively, by a point in Fig. 4. In Figs. 4(a)-(c), we plot the results for (a) $q=1.5$, (b) $q=2$, and (c) $q=3$, where the dashed line denotes the equality case of Eq. (28). In all cases of $q$, the points are above the dashed line, which numerically verifies Eq. (28). The derived relation numerically holds for $q=1.5$, which is non-integer and is a conjectured bound.
Next, we perform numerical calculations for the bounds of the state distribution [Eqs. (31) and (35)]. We generate a random network and consider a continuous time random walk on the network. The transition rate $W_{\mu^{\prime}\mu}$ and the initial node are chosen randomly. We calculate the probability distribution at time $t=\tau$, where $\tau$ is determined randomly, and calculate the Tsallis entropy (Eq. (35)) and its time derivative (Eq. (31)). The range for these random parameters is shown in the caption of Figs. 5 and 6. For each random realization, we plot $\left|\partial_{\tau}H_{q}^{\mathrm{T}}\left[p_{\mu}\left(\tau\mid\mu_{0}\right)\right]\right|$ versus $\sqrt{q\mathcal{A}(\tau)}/(2\tau(q-1))$ (left and right-hand sides of Eq. (31)) in Fig. 5 and $H_{q}^{\mathrm{T}}[p_{\mu}(\tau\mid\mu_{0})]$ versus $(1-e^{-q\tau\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}})/(q-1)$ (left and right-hand sides of Eq. (35)) in Fig. 6 as points. In Figs. 5 and 6, the dashed lines indicate the equality cases in Eqs. (31) and (35), respectively. In Figs. 5 and 6, we plot the results for (a) $q=1.5$, (b) $q=2$, and (c) $q=3$. For all values of $\alpha$, the points lie above the dashed line, providing numerical verification of Eqs. (31) and (35).
Figure 4: Numerical verification of the Tsallis entropy bound for trajectory observables shown in Eq. (28). The Tsallis entropy $H_{q}^{\mathrm{T}}[P(N_{\circ})]$ versus $\left(1-e^{-q\tau\mathfrak{a}(t=0)}\right)/\left(q-1\right)$ is plotted for (a) $q=1.5$, (b) $q=2$, and (c) $q=3$. Random realizations are shown with points and the solid line indicates the equality case of Eq. (28). In (a)-(c), classical simulations begin by selecting the number of states $D$ between 22 and $5$. Then, the transition rate $W_{\mu^{\prime}\mu}$ is randomly produced, and $\tau$ is chosen within the range of $0.1$ to 1010. The observable $N_{\circ}$ is the number of total jumps within $0,\tau]$. The Tsallis entropy of the observable is estimated by averaging the results of $10^{4}$ simulations.
Figure 5: Numerical verification of the time derivative bound of Tsallis entropy of state distribution shown in Eq. (31). The time derivative of Tsallis entropy $\left|\partial_{\tau}H_{q}^{\mathrm{T}}\left[p_{\mu}(\tau)\right]\right|$ versus $\sqrt{q\mathcal{A}(\tau)}/\left(2\tau(q-1)\right)$ is plotted for (a) $q=1.5$, (b) $q=2$, and (c) $q=3$. Random realizations are shown with points and the dashed line indicates the equality case of Eq. (31). In (a)-(c), the simulations begin by selecting the number of states $D$ between 22 and $20$. Then, the transition rate $W_{\mu^{\prime}\mu}$ is randomly produced, and $\tau$ is chosen within the range of $0.1$ to 1010.
Figure 6: Numerical verification of the Tsallis entropy bound of state distribution shown in Eq. (35). The Tsallis entropy $H_{q}^{\mathrm{T}}[p_{\mu}(\tau\mid\mu_{0})]$ versus $\left(1-e^{-q\tau\sum_{\mu\left(\mu\neq\mu_{0}\right)}W_{\mu\mu_{0}}}\right)/\left(q-1\right)$ is plotted for (a) $q=1.5$, (b) $q=2$, and (c) $q=3$. Random realizations are shown with points and the dashed line indicates the equality case of Eq. (35). In (a)-(c), the simulations begin by selecting the number of states $D$ between 22 and $20$. Then, the transition rate $W_{\mu^{\prime}\mu}$ is randomly produced, and $\tau$ is chosen within the range of $0.1$ to 1010.
@@ IV Conclusion
We have introduced replica Markov processes as a general method to derive thermodynamic trade-off relations for nonlinear functionals of probability distributions. By considering $K$ independent identical copies of a Markov process and applying activity-based trade-off relations to suitably chosen replica observables, nonlinear quantities of the original single process can be bounded by linear expectations in the replica process.
Focusing on entropic measures of uncertainty, we have derived upper bounds on the time derivative and the value of the Tsallis entropy for general trajectory-observable distributions in terms of dynamical activity [Eqs. (24) and (25)], as well as Tsallis bounds for general trajectory observable controlled by the initial activity [Eqs. (27) and (28)]. We also have obtained analogous bounds for network diffusion, yielding an upper bound on the time derivative and the value of the Tsallis entropy of the position distribution [Eqs. (31) and (33)] as well as Tsallis entropy bounds for the state distribution controlled by the initial activity [Eqs. (34) and (35)]. In addition, we illustrated that replicas can treat other nonlinear statistics such as extreme-value observables [Eqs. (39)–(41)]. We extended the construction to continuously monitored open quantum systems, where the bounds are expressed via the quantum dynamical activity [Eqs. (51)–(57)].
These results provide entropic counterparts to activity-based uncertainty relations and establish replica processes as a powerful tool for bounding nonlinear information-theoretic quantities in stochastic and quantum thermodynamics.
Appendix A Derivation of Eqs. (24)–(25) and (31)–(33)
@@
$\Lambda_{\bullet}^{(\otimes K)}(\Gamma^{(\otimes K)})\equiv\begin{cases}1&N_{\bullet}^{(1)}(\Gamma^{(1)})=N_{\bullet}^{(2)}(\Gamma^{(2)})=\cdots=N_{\bullet}^{(K)}(\Gamma^{(K)})\\ 0&\mathrm{otherwise}\end{cases}.$ (58)
$\Lambda_{\bullet}^{(\otimes K)}$ returns 1 only when the values of the observable $N_{\bullet}$ in the $K$ processes agree, $N_{\bullet}^{(1)}(\Gamma^{(1)})=N_{\bullet}^{(2)}(\Gamma^{(2)})=\cdots=N_{\bullet}^{(K)}(\Gamma^{(K)})$. Let $\mathcal{Z}_{\bullet}$ be the set of all possible values that the observable $N_{\bullet}(\Gamma)$ can take. The expectation and variance of Eq. (58) are
$\mathbb{E}[\Lambda_{\bullet}^{(\otimes K)}(\Gamma^{\otimes K})]$ $=\sum_{z\in\mathcal{Z}_{\bullet}}P_{\bullet}(z)^{K},$ (59)
$\mathrm{Var}[\Lambda_{\bullet}^{(\otimes K)}(\Gamma^{\otimes K})]$ $=\sum_{z\in\mathcal{Z}_{\bullet}}P_{\bullet}(z)^{K}\left(1-\sum_{z\in\mathcal{Z}_{\bullet}}P_{\bullet}(z)^{K}\right).$ (60)
where we abbreviated $P_{\bullet}(z)\equiv P(N_{\bullet}=z)$. Substituting Eqs. (59) and (60) into Eq. (14), we obtain
$\frac{\left|\partial_{\tau}\sum_{z}P_{\bullet}(z)^{K}\right|}{\sqrt{\sum_{z}P_{\bullet}(z)^{K}\left(1-\sum_{z}P_{\bullet}(z)^{K}\right)}}\leq\frac{\sqrt{K\mathcal{A}(\tau)}}{\tau}.$ (61)
@@
$\left|\frac{d}{d\tau}\arcsin\left(2\sum_{z}P_{\bullet}(z)^{K}-1\right)\right|\leq\frac{\sqrt{K\mathcal{A}(\tau)}}{\tau}.$ (63)
Recall that the following relation holds for $|x|<1$:
$\left|\frac{d}{dx}\left(x\right)\right|\leq\left|\frac{d}{dx}\arcsin\left(x\right)\right|.$ (64)
@@
$\left|\frac{d}{d\tau}\sum_{z}P_{\bullet}(z)^{K}\right|\leq\frac{\sqrt{K\mathcal{A}(\tau)}}{2\tau},$ (65)
@@
Next, we derive Eq. (25). Instead of $N_{\bullet}$, we consider $N_{\circ}$ in Eq. (63). While $N_{\bullet}$ is an arbitrary function of $\Gamma$, $N_{\circ}$ has the additional assumption given in Eq. (6). Let $\mathcal{Z}_{\circ}$ be the set of all possible values that the observable $N_{\circ}(\Gamma)$ can take. By integrating Eq. (63) from $t=0$ to $t=\tau$ and using the triangle inequality, we obtain
@@
By adopting the derivation above, we can obtain the bounds shown in Eqs. (31)–(33). Since $N_{\bullet}(\Gamma)$ can be an arbitrary function of $\Gamma$, we can take
$N_{\bullet}(\Gamma)=X(\tau),$ (68)
@@
which derives Eq. (32) in the main text. When the initial state of the process is fixed to $B_{\mu_{0}}$, indicating $p_{\mu_{0}}(0)=1$, we can derive a tighter inequality than Eq. (70). Specifically, substituting Eq. (68) into Eq. (63), we obtain
@@
Next, we consider the trajectory $\Gamma^{(\otimes K)}$ in the $K$ replica processes. Let $\mathfrak{p}^{(\otimes K)}(\tau)$ be the probability that there is no jump within the interval $0,\tau]$ in the $K$ replica case. Since the replicas are independent processes, from Eq. (77), $\mathfrak{p}^{(\otimes K)}(\tau)$ is bounded from below by
@@
We can relate $\mathfrak{p}^{(\otimes K)}(\tau)$ to the probability $P\left[N^{(\otimes K)}(\Gamma^{(\otimes K)})=0\right]$, which is the probability that the observable $N^{(\otimes K)}$ vanishes. If no jump occurs in the interval $0,\tau]$, then the observable satisfies $N^{(\otimes K)}(\Gamma^{(\otimes K)})=0$ by the condition in Eq. (13). The converse, however, does not necessarily hold. Therefore, the probability that $N^{(\otimes K)}(\Gamma^{(\otimes K)})=0$ is bounded from below by
$P\left[N^{(\otimes K)}(\Gamma^{(\otimes K)})=0\right]\geq\mathfrak{p}^{(\otimes K)}(\tau).$ (79)
@@
where $0<r<s,b\geq 0$ and $b^{r}\leq\mathbb{E}\left[|X|^{r}\right]$ should be satisfied. Using Eq. (80) with $b=0$ for the random variable $N^{(\otimes K)}$, we obtain
$\frac{\mathbb{E}\left[|N^{(\otimes K)}|^{r}\right]^{s/(s-r)}}{\mathbb{E}\left[|N^{(\otimes K)}|^{s}\right]^{r/(s-r)}}$ $\leq P(|N^{(\otimes K)}|>0)$
$=1-P(|N^{(\otimes K)}|=0)$
@@
Appendix C Derivation of Eqs. (27) and (34)
We show the derivation of Eqs. (27) and (34). We consider the particular observable for $N^{(\otimes K)}$ to obtain inequalities for the original single process. Specifically, let us consider the following observable:
$\overline{\Lambda}_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})\equiv\begin{cases}0&N_{\circ}^{(1)}(\Gamma^{(1)})=N_{\circ}^{(2)}(\Gamma^{(2)})=\cdots=N_{\circ}^{(K)}(\Gamma^{(K)})\\ 1&\mathrm{otherwise}\end{cases}$ (82)
$\overline{\Lambda}^{(\otimes K)}_{\circ}$ is similar to Eq. (58). However, the cases in which Eqs. (58) and (82) return 0 and 1 are reversed. $\overline{\Lambda}_{\circ}^{(\otimes K)}$ returns 0 only when the values of the observable $N_{\circ}$ in $K$ processes agree, $N_{\circ}^{(1)}(\Gamma^{(1)})=N_{\circ}^{(2)}(\Gamma^{(2)})=\cdots N_{\circ}^{(K)}(\Gamma^{(K)})$. When there is no jump in all $K$ processes, we trivially obtain $N_{\circ}^{(1)}(\Gamma_{\varnothing}^{(1)})=N_{\circ}^{(2)}(\Gamma_{\varnothing}^{(2)})=\cdots=N_{\circ}^{(K)}(\Gamma_{\varnothing}^{(K)})=0$. Then the observable $\overline{\Lambda}_{\circ}^{(\otimes K)}$ satisfies
@@
which is the condition of Eq. (13). Let $\mathcal{Z}_{\circ}$ be the set of all possible values that the observable $N_{\circ}(\Gamma)$ can take. Then, the expectation becomes the following:

### v3

@@ II.9 Entropic bound on quantum trajectory observable
We aim to quantify the uncertainty associated with quantum trajectory observables. To achieve this, we analyze the probability distribution of $N_{Q}(\Gamma_{Q})$. We derive an upper bound for the Tsallis entropy of this distribution, denoted by $P(N_{Q})$. Following the classical derivation detailed in Appendix A, we obtain
$\left|\frac{\partial}{\partial\tau}H_{q}^{\mathrm{T}}\left[P\left(N_{Q}\right)\right]\right|\leq\frac{\sqrt{q\mathcal{B}(\tau)}}{2\tau(q-1)},$ (56)
where $q=2,3,\cdots$. Equation (56) holds for an arbitrary observable $N_{Q}(\Gamma_{Q})$ of the quantum trajectory $\Gamma_{Q}$. Again, for $N_{Q\circ}$, which additionally satisfies the condition of Eq. (48), the following relation holds:
@@
where $q=2,3,\cdots$, and the following condition should be satisfied:
@@
Equations (56) and (57) are quantum generalizations of Eqs. (24) and (25), respectively, and thus their derivations are the same as those of Eqs. (24) and (25). From Eq. (54), we obtain the following relation (see Appendix F for details):
@@
where $q=2,3,\cdots$. Equation (59) is a quantum generalization of Eq. (28). Equations (57) and (59) are similar in that the left-hand side is identical and the right-hand sides comprise the quantum dynamical activity. It can be shown that Eq. (59) provides a tighter bound. In the classical case, the left-hand side of Eq. (28) is the Tsallis entropy of the observable $N_{\circ}$, and the right-hand side depends on the dynamical activity $\mathfrak{a}(t=0)$ at time $t=0$. On the other hand, in the quantum case, the right-hand side of Eq. (59) depends on the time integral of the quantum dynamical activity $\mathcal{B}(t)$.
Figure 3: Description of networks. (a) Illustration of the dependence of the entropic bounds for the state distribution [Eqs. (35) and (36)]. The entropic bounds shown in Eqs. (35) and (36) depend on $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$, where $B_{\mu_{0}}$ is the initial state, i.e., the local escape rate from $B_{\mu_{0}}$. The local escape rate from $B_{\mu_{0}}$ is the sum of transition rates to the nodes reachable from the initial state in one jump. Suppose the initial state is $B_{1}$, which is colored dark red. Then, the reachable states are represented by purple nodes. Therefore, the local escape rate becomes $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}=W_{21}+W_{31}+W_{41}$. (b) Twitter interaction network used in our numerical simulations. The data [72, 73], available from [74, 75], represent interactions among members of the 117th United States Congress as a directed, weighted graph. Edge weights are empirically estimated transmission probabilities. The network contains 475475 nodes and 13,28913{,}289 edges.
@@ III Numerical simulation
We run numerical simulations to verify the derived bounds. We use Twitter interaction network data [72, 73], which are available from [74, 75]. This dataset captures interactions among members of the 117th United States Congress as a directed, weighted social network. The overall network configuration is shown in Fig. 3. Each edge weight represents an empirically estimated transmission probability from one member to another. This probability was computed by counting how often a member’s tweets are retweeted, quote-tweeted, replied to, or mentioned by another member and then normalizing that total by the number of tweets produced by the potential influencer. This network is well suited for studying information diffusion and influence dynamics in social systems. In total, the number of nodes is $475$, and the number of edges is $13,289$. We construct a continuous-time Markov process from the dataset, assigning transition rates based on the transmission probabilities. Because these probabilities are very small, we approximate the transition rates by using the corresponding transition probabilities. We analyze the degree of information diffusion by computing the generalized entropies $H_{q}^{\mathrm{T}}$ or $H_{\alpha}^{\mathrm{R}}$ and their upper bounds.
We begin by examining the bound in Eq. (31). To quantify how information spreads through the network, we use the Tsallis entropy $H_{q}^{\mathrm{T}}[p_{\mu}(\tau)]$. Our focus is on the diffusion speed, measured by $\left|\partial_{\tau}H_{q}^{\mathrm{T}}\left[p_{\mu}\left(\tau\right)\right]\right|$. For this calculation, we assume that the initial state is the node with the maximum out-degree in the network. In Fig. 4, we plot $\left|\partial_{\tau}H_{q}^{\mathrm{T}}\left[p_{\mu}\left(\tau\right)\right]\right|$ as a function of $\tau$ for (a) $q=2$ and (b) $q=3$, where the dashed line denotes the upper bound $\sqrt{q\mathcal{A}(\tau)}/(2\tau(q-1))$ given on the right-hand side of Eq. (31). From Fig. 4, it can be seen that, for both $q=2$ and $q=3$, the derivative values of the Tsallis entropy fall below the upper bound, confirming that Eq. (31) holds. The derivative value of the Tsallis entropy is larger in the region where $\tau$ is small. This indicates that diffusion occurs more rapidly in the early stages, when $\tau$ is small. Furthermore, in both cases, a gap from the upper bound is observed in the region where $\tau$ is small.
Next, we examine the entropy at time $\tau$ and its upper bound. Here, we use the Rényi entropy [Eq. (21)] rather than the Tsallis entropy to quantify the degree of diffusion. The reason for employing the Rényi entropy is that, in the upper bound of the Tsallis entropy [Eq. (36)], the dynamical activity and $\tau$ are included in the exponent, which causes the upper bound to rapidly converge to 1/(q1)1/(q-1). Again, we assume that the initial state is the node with the maximum out-degree. In Fig. 5, we plot $H_{\alpha}^{\mathrm{R}}[p_{\mu}(\tau\mid\mu_{0})]$ as a function of $\tau$ for (a) $\alpha=2$ and (b) $\alpha=3$, where the dashed line denotes the upper bound $\alpha\tau/(\alpha-1)\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$ given on the right-hand side of Eq. (35). It can be seen that the bound represented by Eq. (35) provides a very tight upper bound up to approximately $\tau=2$. In the region where $\tau$ is greater than $3$, while the entropy value is nearly saturated, the upper bound increases linearly with respect to $\tau$, resulting in a gap between the upper bound and the entropy value. The behaviors for $\alpha=2$ and $\alpha=3$ are similar, but around the region where the slope of the entropy changes, $\alpha=3$ changes more steeply.
Finally, we compare the left and right sides of Eq. (35) for different initial settings. We calculate $H_{\alpha}^{\mathrm{R}}[p_{\mu}(\tau)]$ and $\alpha\tau/(\alpha-1)\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$, which are the left and right sides of Eq. (35) at time $\tau=10$. There are $475$ nodes in the network, and using all of these nodes as initial states, the results for each of the $475$ configurations are plotted as points on Fig. 6 for (a) $\alpha=2$ and (b) $\alpha=3$. The equality case of Eq. (35) is shown by the dashed lines in Fig. 6. Because all the points in Figs. 6(a) and (b) are above the dashed lines, we can numerically validate Eq. (35).
Figure 4: Numerical verification of the time derivative of the Tsallis entropy bound for the state distribution shown in Eq. (31) using the Twitter interaction network. The time derivative of the Tsallis entropy, $\left|\partial_{\tau}H_{q}^{\mathrm{T}}\left[p_{\mu}\left(\tau\right)\right]\right|$, is plotted as a function of $\tau$ for (a) $q=2$ and (b) $q=3$ with solid lines. The dashed lines denote the upper bound in Eq. (31), which is (a) $\sqrt{2\mathcal{A}(\tau)}/(2\tau)$ and (b) $\sqrt{3\mathcal{A}(\tau)}/(4\tau)$. The initial state is $p_{\mu_{0}}(0)=1$, where $B_{\mu_{0}}$ is the node with the maximum out-degree in the Twitter interaction network.
Figure 5: Numerical verification of the Rényi entropy bound for the state distribution shown in Eq. (35) using the Twitter interaction network. The Rényi entropy $H_{\alpha}^{\mathrm{R}}[p_{\mu}(\tau\mid\mu_{0})]$ is plotted as a function of $\tau$ for (a) $\alpha=2$ and (b) $\alpha=3$ using solid lines. The dashed lines denote the upper bounds in Eq. (35), which are (a) $2\tau\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$ and (b) $(3\tau/2)\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$. The initial state is $p_{\mu_{0}}(0)=1$, where $B_{\mu_{0}}$ is the node with the maximum out-degree in the Twitter interaction network.
Figure 6: Numerical verification of the Rényi entropy bound for the state distribution shown in Eq. (35) using the Twitter interaction network. The Rényi entropy $H_{\alpha}^{\mathrm{R}}[p_{\mu}(\tau\mid\mu_{0})]$ versus $\alpha\tau/(\alpha-1)\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$ at time $\tau=10$ is plotted for (a) $\alpha=2$ and (b) $\alpha=3$. Each point corresponds to a different initial state, and we use every 475475 node as the initial state. The dashed lines denote the equality case in Eq. (35).
@@ IV Conclusion
We have introduced replica Markov processes, collections of $K$ independent, identical copies of a Markov process, as a systematic framework for deriving thermodynamic trade-off relations that constrain nonlinear functionals of probability distributions. By applying existing activity-based trade-off relations [Eqs. (14) and (17)] to suitably chosen observables in the replica process, nonlinear quantities of the original single process are expressed as linear expectations in the enlarged state space. The replicas serve as purely virtual constructs, analogous to multi-copy measurements in quantum information and the replica trick in spin-glass theory, so that the final results are statements about the single process alone. Focusing on entropic measures of uncertainty, we have obtained several families of bounds:
  1. 1.
    Trajectory-observable entropies. Upper bounds on the time derivative [Eq. (24)] and the value [Eq. (25)] of the Tsallis entropy for the distribution of a general trajectory observable, expressed in terms of the time-integrated dynamical activity $\mathcal{A}(\tau)$. Additionally, Rényi and Tsallis entropy bounds [Eqs. (27) and (28)] controlled by the initial dynamical activity $\mathfrak{a}(t{=}0)$ have been derived.
  2. 2.
    State-distribution entropies and network diffusion. Analogous bounds on the time derivative [Eq. (31)] and the value [Eq. (34)] of the Tsallis entropy for the state probability distribution $p_{\mu}(\tau)$, which quantifies the degree of diffusion on a network. Notably, the Rényi and Tsallis entropy bounds [Eqs. (35) and (36)] depend only on the observation time $\tau$ and the local escape rate from the initial node, $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$, irrespective of the global network size or topology.
  3. 3.
    Extreme-value observables. We have demonstrated that the replica construction naturally extends to other nonlinear statistics. In particular, trade-off relations for the maximum of $M$ independent realizations [Eqs. (39)–(43)] have been obtained, showing that the entropic uncertainty of extreme-value observables is similarly governed by the dynamical activity.
  4. 4.
    Quantum extension. The entire framework has been generalized to continuously monitored open quantum systems described by Lindblad dynamics. The resulting bounds [Eqs. (56), (57), and (59)] take the same structural form as their classical counterparts, with the classical dynamical activity replaced by the quantum dynamical activity $\mathcal{B}(\tau)$, which additionally accounts for coherent contributions to the dynamics.
Numerical simulations on a real-world Twitter interaction network among members of the 117th United States Congress have confirmed the validity and tightness of the derived bounds, particularly the Rényi entropy bound for the state distribution [Eq. (35)]. Our framework extends the scope of activity-based trade-off relations from linear functionals of probability distributions to genuinely nonlinear information-theoretic quantities, which are inaccessible to conventional methods.
Appendix A Derivation of Eqs. (24)–(25)
@@
$\Lambda^{(\otimes K)}(\Gamma^{(\otimes K)})$
$\equiv\begin{cases}1&N^{(1)}(\Gamma^{(1)})=N^{(2)}(\Gamma^{(2)})=\cdots=N^{(K)}(\Gamma^{(K)})\\ 0&\mathrm{otherwise}\end{cases}.$ (60)
$\Lambda^{(\otimes K)}$ returns 1 only when the values of the observable $N$ in the $K$ processes agree, $N^{(1)}(\Gamma^{(1)})=N^{(2)}(\Gamma^{(2)})=\cdots=N^{(K)}(\Gamma^{(K)})$. Let $\mathcal{Z}$ be the set of all possible values that the observable $N(\Gamma)$ can take. The expectation and variance of Eq. (60) are
$\mathbb{E}[\Lambda^{(\otimes K)}(\Gamma^{(\otimes K)})]$ $=\sum_{z\in\mathcal{Z}}P(z)^{K},$ (61)
$\mathrm{Var}[\Lambda^{(\otimes K)}(\Gamma^{(\otimes K)})]$ $=\sum_{z\in\mathcal{Z}}P(z)^{K}\left(1-\sum_{z\in\mathcal{Z}}P(z)^{K}\right),$ (62)
where we abbreviate $P(z)\equiv P(N(\Gamma)=z)$. Substituting Eqs. (61) and (62) into Eq. (14), we obtain
$\frac{\left|\partial_{\tau}\sum_{z}P(z)^{K}\right|}{\sqrt{\sum_{z}P(z)^{K}\left(1-\sum_{z}P(z)^{K}\right)}}\leq\frac{\sqrt{K\mathcal{A}(\tau)}}{\tau}.$ (63)
@@
$\left|\frac{d}{d\tau}\arcsin\left(2\sum_{z\in\mathcal{Z}}P(z)^{K}-1\right)\right|\leq\frac{\sqrt{K\mathcal{A}(\tau)}}{\tau}.$ (65)
Recall that the following relation holds for $|f(\tau)|<1$:
$\left|\frac{d}{d\tau}f(\tau)\right|\leq\frac{1}{2}\left|\frac{d}{d\tau}\arcsin(2f(\tau)-1)\right|.$ (66)
@@
$\left|\frac{d}{d\tau}\sum_{z\in\mathcal{Z}}P(z)^{K}\right|\leq\frac{\sqrt{K\mathcal{A}(\tau)}}{2\tau},$ (67)
@@
Next, we derive Eq. (25). Instead of $N(\Gamma)$, we consider $N_{\circ}(\Gamma)$ in Eq. (65). While $N(\Gamma)$ is an arbitrary function of $\Gamma$, $N_{\circ}(\Gamma)$ has the additional assumption given in Eq. (5). Let $\mathcal{Z}_{\circ}$ be the set of all possible values that the observable $N_{\circ}(\Gamma)$ can take. By integrating Eq. (65) from $t=0$ to $t=\tau$ and using the triangle inequality, we obtain
@@
Appendix B Derivation of Eqs. (31), (32), and (34)
By adopting the derivation above, we can obtain the bounds shown in Eqs. (31), (32), and (34). Since $N(\Gamma)$ can be an arbitrary function of $\Gamma$, we can take
$N(\Gamma)=X(\tau),$ (70)
@@
which derives Eq. (32) in the main text. When the initial state of the process is fixed to $B_{\mu_{0}}$, implying $p_{\mu_{0}}(0)=1$, we can derive a tighter inequality than Eq. (72). Specifically, substituting Eq. (70) into Eq. (65), we obtain
@@
Next, we consider the trajectory $\Gamma^{(\otimes K)}$ in the $K$ replica processes. Let $\mathfrak{p}^{(\otimes K)}(\tau)$ be the probability that there is no jump within the interval $0,\tau]$ in the $K$-replica case. Since the replicas are independent processes, from Eq. (79), $\mathfrak{p}^{(\otimes K)}(\tau)$ is bounded from below by
@@
We can relate $\mathfrak{p}^{(\otimes K)}(\tau)$ to the probability $P\left[N_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})=0\right]$, which is the probability that the observable $N_{\circ}^{(\otimes K)}$ vanishes. If no jump occurs in the interval $0,\tau]$, then the observable satisfies $N_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})=0$ by the condition in Eq. (13). The converse, however, does not necessarily hold. Therefore, the probability that $N_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})=0$ is bounded from below by
$P\left[N_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})=0\right]\geq\mathfrak{p}^{(\otimes K)}(\tau).$ (81)
@@
where $0<r<s$, $b\geq 0$, and $b^{r}\leq\mathbb{E}\left[|X|^{r}\right]$ should be satisfied. Using Eq. (82) with $b=0$ for the random variable $N_{\circ}^{(\otimes K)}$, we obtain
$\frac{\mathbb{E}\left[|N_{\circ}^{(\otimes K)}|^{r}\right]^{s/(s-r)}}{\mathbb{E}\left[|N_{\circ}^{(\otimes K)}|^{s}\right]^{r/(s-r)}}$ $\leq P(|N_{\circ}^{(\otimes K)}|>0)$
$=1-P(|N_{\circ}^{(\otimes K)}|=0)$
@@
Appendix D Derivation of Eqs. (27) and (35)
We show the derivation of Eqs. (27) and (35). We consider a particular observable to obtain inequalities for the original single process. Specifically, let us consider the following observable:
$\overline{\Lambda}_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})$
$\equiv\begin{cases}0&N_{\circ}^{(1)}(\Gamma^{(1)})=N_{\circ}^{(2)}(\Gamma^{(2)})=\cdots=N_{\circ}^{(K)}(\Gamma^{(K)})\\ 1&\mathrm{otherwise}\end{cases}$ (84)
$\overline{\Lambda}^{(\otimes K)}_{\circ}$ is similar to Eq. (60). However, the cases in which Eqs. (60) and (84) return 0 and 1 are reversed. $\overline{\Lambda}_{\circ}^{(\otimes K)}$ returns 0 only when the values of the observable $N_{\circ}$ in the $K$ processes agree, $N_{\circ}^{(1)}(\Gamma^{(1)})=N_{\circ}^{(2)}(\Gamma^{(2)})=\cdots N_{\circ}^{(K)}(\Gamma^{(K)})$. When there is no jump in all $K$ processes, we trivially obtain $N_{\circ}^{(1)}(\Gamma_{\varnothing}^{(1)})=N_{\circ}^{(2)}(\Gamma_{\varnothing}^{(2)})=\cdots=N_{\circ}^{(K)}(\Gamma_{\varnothing}^{(K)})=0$. Then the observable $\overline{\Lambda}_{\circ}^{(\otimes K)}$ satisfies
@@
which is the condition of Eq. (13). Then, the expectation becomes the following: